OpenStudy (anonymous):

can someone solve and show steps for (x+iy)^2=18i

myininaya (myininaya):

multiply (x+iy)(x+iy) write what you got when you do that

OpenStudy (anonymous):

solve the radical equations and remember to weed out extraneous solutions. \[\sqrt{2x^2+2}=3-x\]

myininaya (myininaya):

square both sides 2x^2+2=(3-x)^2 2x^2+2=9-6x+x^2 now solve for x and check both solutions